OpenStudy (anonymous):

Jason had $32. He spent all the money on buying 4 CDs for $x each and 2 magazines for $y each. If Jason would have bought 5 CDs and 2 magazines, he would have run short of $4. A student concluded that the price of each CD is $5 and the price of each magazine is $6. Which statement best justifies whether the student's conclusion is correct or incorrect?

OpenStudy (anonymous):

ANSWER CHOICES The student's conclusion is incorrect because the solution to the system of equations 4x + 2y = 4 and 5x + 2y = 36 is (9, 4). The student's conclusion is incorrect because the solution to the system of equations 4x + 2y = 32 and 5x + 2y = 36 is (4, 8). The student's conclusion is correct because the solution to the system of equations 4x + 2y = 4 and 5x + 2y = 36 is (5, 6). The student's conclusion is correct because the solution to the system of equations 4x + 2y = 32 and 5x + 2y = 36 is (5, 6).

OpenStudy (jewlzme17):

The answer would be B

OpenStudy (anonymous):

@jewlzme17 is it just your guess?

OpenStudy (jewlzme17):

no, that's the answer..i did the problem. I am really good at math!

OpenStudy (anonymous):

@jewlzme17 ok, can you help me with other questions then too ?

OpenStudy (jewlzme17):

ya :) what level of math are you in?

OpenStudy (anonymous):

algebra 1 honors

OpenStudy (anonymous):

1/2x + 3/2=2^x what is the solution x = 1 x = 2 x = 4 x = 8

OpenStudy (anonymous):

ahh.. i'll buy it at a high price!

OpenStudy (jewlzme17):

ya just one sec, um what exactly is this under? like the topic of it all because i want to make sure its correct and i'm not using the incorrect form because there are a few ways you could do this but then you get diff answers.

OpenStudy (anonymous):

its in the chapter of system of equations

OpenStudy (anonymous):

@jewlzme17 can you help or not? i dont have much time left at all

OpenStudy (jewlzme17):

and there is no "Y"? that's odd since it is in the chapter system of equations but I got "1" but you might want to check that because i'm not positive because it was a weird setup.

OpenStudy (anonymous):

yeah its so weird, but im gonna go with 1, thanks, can you help me with more ?

OpenStudy (jewlzme17):

no problem aha sorry if its wrong but sure

OpenStudy (anonymous):

its ok, at least you tried, The tables below show the values of f(x) and g(x) for different values of x. f(x) = 2(3)x Based on the tables, what is the solution to the equation 2(3)x = 3x + 3?

OpenStudy (anonymous):

answer choices : x = –1 x = 0 x = 1 x = 6

OpenStudy (sleepyjess):

@jewlzme17 ... you need to explain the problem before giving the answer... just giving the answer doesn't help anyone

OpenStudy (sleepyjess):

@yeah_i_suck_at_math ... what do you mean you don't have much time left?

OpenStudy (the_fizicx99):

What did you mean by "It helps me to be honest"

OpenStudy (sleepyjess):

Giving just the answers is a violation of the rules of this site.

OpenStudy (sleepyjess):

And since you are almost out of time, does that mean you are taking a test?

OpenStudy (jewlzme17):

what exactly is that chart showing? like do you just need to fill in anything or can you just solve the equation just like that?

OpenStudy (the_fizicx99):

Why're you erasing your replies lol

OpenStudy (sleepyjess):

We aren't allowed to help you on a test.

OpenStudy (the_fizicx99):

Curious.

OpenStudy (anonymous):

@jewlzme17 i dont know, i thought it was just one of those sequence functions things, im confused

OpenStudy (sleepyjess):

Do you realize that teachers are on here also? And that if you are caught cheating on a test, you will get an F?

OpenStudy (the_fizicx99):

What?! Wtf does that even mean?!

OpenStudy (the_fizicx99):

-_-

OpenStudy (jewlzme17):

okay i'm sorry that i gave her an answer..i do not want to get into drama and into all this violation stuff so all you need to do for this is just follow everything the chart tells you and fill it in and you will get your answer! ahah bye!

OpenStudy (anonymous):

@jewlzme17 my most educated guess was x=1 but idk ...

OpenStudy (jewlzme17):

well yes when you solve 2(3)x = 3x + 3 point blank then you will get 1! and thank you ahah and it just because it happened to my brother and he got kicked off this whole thing for helping someone out and like he wasn't just giving answers but someone didn't like him and reported him for just giving answers and he got kicked off so ya i don't want it to happen to me

OpenStudy (the_fizicx99):

xD or until your IP finally gets banned. :) They can see your IP and decide if they can ban it. That's if your not in a library, otherwise its a possibility.

OpenStudy (the_fizicx99):

FYI, posting exam questions will also get you suspended.

OpenStudy (anonymous):

oh, the tragedy!